### v1

@@ 7 Conclusion
In this work, we analyzed the inherent susceptibility of stream-based anonymous communication to traffic analysis attacks. In particular, we focused on intersection attacks because these constitute the class of attacks which streams are fundamentally vulnerable against, compared to independent messages. We introduced progressive pruning, a methodology to quantify the anonymity level of streams, with regard to the effectiveness of an intersection attack. With that, we contribute a new perspective: Instead of regarding the susceptibility to intersection attacks as a binary property, we enable a more differentiated view. The ability to compare the anonymity achieved by different system variants can be used for designing new ACNs, taking attackers’ cost into account.
We carried out simulations to investigate the impact of different stream properties in tailored synthetic scenarios. Moreover, we demonstrated how progressive pruning can be applied to large systems like the Tor network, estimating anonymity factors with regard to network topology. Our results indicate that the susceptibility towards traffic analysis attacks of streams is primarily dependent on the length of streams, relative to each other. Longer streams suffer from worse anonymity because they can more easily be distinguished from others by considering their number of messages. At the same time, they serve as cover traffic, increasing anonymity for shorter streams. When looking at Tor as an example of stream-based anonymity networks, we observe different influences when scaling the network. On the one hand, scaling the network horizontally (adding relays) helps increase anonymity set sizes in the context of our anonymity estimate that takes the perspective of a global adversary who monitors incoming and outgoing traffic of the anonymity system. On the other hand, when considering a vertically scaled network, we can see that an increased number of users is beneficial for anonymity.
We argue that stream-based communication is of high practical relevance for online privacy and should therefore be a goal when designing future anonymity systems. While our work shows that it has inherent, negative consequences for anonymity due to being vulnerable to intersection attacks, our results may serve as a starting point in working around these limitations with the vision of combining stream-based communication with strong anonymity.

### v2

@@ 6 Conclusion
In this work, we analyzed the inherent vulnerability of stream-based anonymous communication to traffic analysis. We specifically focused on intersection attacks, which exploit fundamental weaknesses of streams compared to independent messages. To this end, we introduced progressive pruning, a methodology to quantify stream anonymity in the face of such attacks, enabling a differentiated assessment beyond a binary view of susceptibility. This approach facilitates comparing anonymity across system variants, aiding the design of new ACNs that account for attackers’ costs.
We conducted simulations to examine the impact of stream properties in various scenarios. Our findings highlight that stream length relative to others significantly influences susceptibility: longer streams are more easily distinguished, reducing their anonymity, but also serve as cover traffic for shorter streams.
Additionally, we demonstrated the application of progressive pruning to large real-world networks, By using Tor as a case study, we observed that horizontal scaling improves anonymity by increasing anonymity set sizes under a global adversary model, but anonymity is distributed very unevenly. On the other hand, when considering a vertically scaled network, we can see that an increased number of users is beneficial for anonymity.
Despite the inherent vulnerability to intersection attacks, our findings provide a methodology to systematically explore mitigation strategies increasing the difficulty faced by adversaries.